OpenStudy (anonymous):

Complete the square x^2 + 2x - 3 = 0

OpenStudy (anonymous):

x^2+2x-3=0 +4 to each side x^2+2x+1=4 (x+1)^2=4 x+1=2 or x+1=-2 x=1 or x=-3

OpenStudy (anonymous):

Thanks sooooo much :)

OpenStudy (anonymous):

Remember the general steps for completing the square is to first make it so that the constant on the left is equal to the square of half of the coefficient of the x term. Then it factors into x+/- the sqrt of the constant.